solve for x using the addition principle: - 8 + x = -24
solve for y using the addition principle: - 1/8 + y = - 3/4
solve for a using the addition and multiplication principles: 14 - 6a = -2a + 3

Answer by stanbon(75887)

solve for x using the addition principle: - 8 + x = -24
Add 8 to both sides to get:
x = -16
-------------
solve for y using the addition principle: - 1/8 + y = - 3/4
Add 1/8 to both sides to get:
y = -6/8 + 1/8
y = -5/8
-----------------------
solve for a using the addition and multiplication principles:
14 - 6a = -2a + 3
Add 6a to both sides to get:
14 = 4a + 3
Subtract 3 from both sides to get:
4a = 11
Divide both sides by 4 to get:
a = 11/4
------------------------
Cheers,
Stan H.